OpenStudy (heyitszia1):

PLEASE HELP ME!! FAN AND MEDAL The table below shows the ticket rates for whale watching trips offered by Dolphin Tricks & Tours: Age Ticket Price Under 3 years free 3 to 12 years $35 Over 12 years $46 On a certain day, the company took 88 people on whale watching trips. There were 44 children aged 12 and under, of which some children were under 3 years. If x represents the number of children under 3 years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day?

OpenStudy (heyitszia1):

(44 + x)35 + (88 − (44 + x)) ⋅ 46 = C (88 − x)35 + (88 − 44) ⋅ 46 = C (44 − x)35 + (88 − 44) ⋅ 46 = C (88 + x)35 + (88 − (44 + x)) ⋅ 46 = C